my husband is four time older then my daughter. in twenty years time, he will be twice as old as my daughter. how old is my daughter?
Answer by mananth(16949)

Daughter x
Dad 4x
20 years later
Daughter =x+20
Dad=4x+ 20
Dad=2 times older
4x+20=2 (x+20)
4x+20=2 x+40
2x=40-20
x=10
x=10 Daughter's age now